# Constraining the Nuclear Equation of State of neutron star via high-frequency quasi-periodic oscillation in short gamma-ray bursts

### v1

@@ 3 Radial oscillation
The first direct evidence of short GRB origin, which comes from the merger of NS–NS is the detection of GW 170817 by Advanced LIGO and Virgo associated with short GRB 170817A and an optical/infrared transient known as kilonova AT2107gfo (Abbott et al., 2017; Goldstein et al., 2017; Zhang et al., 2018). From a theoretical point of view, there are four different types of remnants of binary NS mergers, such as a black hole, a hypermassive NS, a supramassive NS supported by rigid rotation with a lifetime of hundreds of seconds before collapsing into a BH, and a stable NS (Dai & Lu, 1998a, b; Dai et al., 2006; Rosswog et al., 2000; Fan & Xu, 2006; Rezzolla et al., 2010; Giacomazzo & Perna, 2013; Zhang, 2013; Lasky et al., 2014; Lü et al., 2017; Gao et al., 2017). If the remnant is a newborn NS, the temperature is expected to reach several MeV, or even exceed tens of MeV in the immediate aftermath of the merger (Oechslin et al., 2007; Bauswein et al., 2010). Such high temperatures already exceed the point at which a crystalline lattice is stable, and the initial state of the remnant from NS-NS should be fluid-filled. The cooling timescale is expected to range from thousands of seconds to days before the crystal forms.. Therefore, torsional oscillations, which are shear forces from a crystalline crust, are unlikely to occur, but radial modes are more suitable (Stergioulas et al., 2011). The radial oscillation is driven by pressure and gravity, producing a radial displacement that is directly affected by the density and pressure profiles of NS (Chandrasekhar, 1964; Chanmugam, 1977; Gondek et al., 1997a). Gondek et al. (1997b) found that the frequency spectrum of the lowest modes differs significantly from that of cold NS. In this section, we will discuss the radial oscillation of NS and constrain the EOS of NS by adopting the high-frequency QPO of short-duration GRB 931103B and GRB 910711, assuming these two short GRBs are indeed from the merger of NS-NS.
@@ 3.1 Physical model
Within the framework of general relativity, the study of infinitesimal radial adiabatic oscillations in stars was initially proposed by Chandrasekhar (1964), and applied to enhance its utility for numerical calculation by Chanmugam (1977) later. The study focused on two essential quantities: one being the radial displacement ($\Delta r$) which represents the radial motion of a matter element relative to its equilibrium position. The other being the Lagrangian pressure perturbation ($\Delta P$), which quantifies the change in pressure associated with the displacement. These quantities are determined by a pair of ordinary differential equations, which are expressed as follows:
@@
Here, $\xi=\Delta r/r$ is the normalized radial displacement. To avoid the singularity in $\xi$ at the stellar center and ensure a well-behaved normalization for the eigen-functions, we adopt $\xi(0)=1$ as the boundary condition. $\omega_{n}=2\pi\nu_{n}$ is the angular frequency of radial oscillation with the eigen-frequency $\nu_{n}$, the stability condition of oscillations requires $\omega^{2}\geq 0$. $\Gamma$ is the adiabatic index which is related to the baryon number density $n_{b}$ and pressure (Baym et al., 1971), namely,
@@
By solving Eqs. (7) and (8), it is necessary to apply boundary conditions at the stellar center and surface, namely,
@@
If this is the case, Eqs. (7) and (8) can be reduced to a second-order differential equation for $\xi(r)$, and switched to the form of a Sturm-Liouville eigenvalue problem. Similar to the case of torsional oscillation, the eigenvalues $\omega^{2}$ can also yield a discrete spectrum $\omega_{0}^{2}<\omega_{1}^{2}<\dots<\omega_{n}^{2}$, and each eigenvalue corresponds to an eigenfunction $\xi_{n}$ that has $n$ nodes within $0\leq r\leq R$. The fundamental mode ($n=0$), also known as the f-mode, corresponds to the lowest eigenvalue, while the higher modes with $n=1,2,3,\dots$ are referred to as p-modes (Sagun et al., 2020). The eigen-frequencies not only describe the nature of stellar oscillations but also serve as critical indicators of the system’s stability under small perturbations. Hence, all eigenvalues should satisfy $\omega_{n}^{2}\geq 0$ for every $n$, and ensure that a static stellar model is dynamically stable under small, radial, and adiabatic perturbations.
@@ 3.2 Constraint on the EOS of NS
HMNS, supported by differential rotation which survives 10–100 ms before collapsing into a BH, is a possible remnant of a binary NS merger. Initially, the HMNS is expected to have an extremely high temperature, potentially reaching tens of MeV. Such high temperatures significantly affect the physical properties of HMNS, such as the EOS of NSs. By assuming that the high-frequency QPOs in the short-duration GRB 931101B and GRB 910711 are caused by the radial oscillation of HMNS from a binary NS merger, we employ six EOSs for NSs (NL3, IUF, TM1, TMA, FSG, and BHBLp), which are suitable for the condition of high temperature. The interpolated data for the selected EOS models were obtained from the publicly available CompOSE database [^1] https://compose.obspm.fr/table. (Oertel et al., 2017; Typel et al., 2015).
Figs. 3 and 4 show the eigen-frequencies ($\nu_{n}$) of various EOS models as functions of the central density ($\rm log\rho_{c}$) with different temperatures for GRB 931101B and GRB 910711, respectively. The maximum temperature of five EOSs (i.e., IUF, TM1, TMA, FSG and BHBLp) can reach 40 MeV, while for NL3, it extends up to as high as 60 MeV. It is found that the frequencies of $n=1$ modes at zero temperature are significantly higher than those of observed values for those two short GRBs. However, the frequency of $n=1$ mode is rapidly decreased when the temperature is increased. The shaded vertical regions in Fig. 3 and Fig. 4 correspond to the ranges of $\rho_{c}$ where the eigen-frequencies of $n=0$ and $n=1$ modes match the observed lower- and higher-frequencies of QPO, respectively. If this is the case, the temperature at which the observed lower- and higher-frequencies of the QPO correspond to the same $\rho_{c}$ needs to be determined and can be used to constrain the EOS of NS.
To find the temperature at which the vertical bands overlap, we define a metric: $\text{Metric}=\rho_{\rm gap}-\rho_{\rm cross}$, which quantifies the spatial relationship between the two vertical bands. Here, the $\rho_{\rm gap}$ represents the distance between the closest edges of the two regions, namely:
@@
$[x_{0,\text{min}},x_{0,\text{max}}]$ and $[x_{1,\text{min}},x_{1,\text{max}}]$ are the $\rho_{c}$ ranges obtained by comparing the $n=0$ mode with the observed lower-frequency QPO and the $n=1$ mode with the higher-frequency QPO, respectively. If the bands overlap, $\rho_{\rm gap}=0$. $\rho_{\rm cross}$ is defined as
@@
If the bands do not overlap, $\rho_{\rm cross}<0$; otherwise, it is the size of the overlapping region between the two bands. $\text{Metric}=0$ means that the temperature at which the observed lower and higher frequencies of the QPO correspond to the same $\rho_{c}$. Fig. 5 shows the relationship between the Metric and temperature ($T$) for different EOS. It is found that most EOS allow the Metric to reach a negative value at a specific temperature for those two short GRBs, except for the stiffest EOS, NL3. So that, except NL3, we cannot rule out the other five EOS based on the observed lower- and higher-frequencies of QPO for those two short GRBs.
@@
On the other hand, the effect of redshift in the above calculation is ignored due to non-redshift measured. In fact, one needs to consider the uncertainty caused by the redshift of these two short GRBs. Fig. 6 shows the relationship between the temperature $T$ and the redshift $z$ for different EOSs. Here, temperature $T$ values are corresponding to the points where the Metric reaches its minimum value. For example, the minimum value of temperature in Fig. 5 for a given EOS is corresponding to $z=0$. Then, we perform the calculations by varying the redshift range from 0 to 1 based on the observations [^2] Lü et al. (2015) found that the redshift distribution for short GRBs with redshift measured is from 0 to 1, and average redshift is about $z=0.58$.. We find that the temperature decreases as the redshift increases, and it is due to the relativistic effect of redshift which increases the frequency in the rest frame. Moreover, it is found that the NL3 EOS can reach the overlap of $\rho_{c}$ ranges at a higher temperature compared to other EOSs when $z\gtrsim 0.5$. For example, Fig. 7 shows the eigen-frequency as a function of the NS mass for $n=0$ (solid lines) and $n=1$ (dashed lines) oscillation modes for different EOSs at $z=0$ and $z=1$, respectively.
@@
In any case, it is difficult to constrain the EOS of NS based solely on the observed high-frequency QPO in those two short GRBs, and cannot rule out any EOS by considering the effect from both temperature and redshift. However, it is found that the different EOS of NS correspond to different temperatures for a given redshift with $\text{Metric}=0$. In other words, if we believe that the observed high-frequency QPOs in short GRBs are indeed caused by radial oscillations of HMNS from binary NS mergers, one can constrain the EOS only when the redshift and temperature of the remnant can be measured.
@@ 4 Conclusion and Discussion
Neutron stars offer unparalleled opportunities for investigating extreme physics phenomena in the universe. Gaining insights into the composition of matter and the structure of NSs would be a significant advancement in nuclear physics and astrophysics. The application of seismology in QPOs of SGRs has demonstrated immense potential in this regard, particularly in providing constraints on the EOS and the thickness of the crust of NSs. The recent identification of QPOs from two short-duration GRBs (931101B and 910711) may provide a new opportunity to explore the nature of NSs by employing seismology.
@@
In this paper, we investigate the potential of torsional and radial oscillation models to explain the observed high-frequency QPO in two short-duration GRBs (931101B and GRB 910711), and then try to constrain the EOS of NS. The following interesting results can be obtained.
  • $\bullet$
@@
    Torsional oscillation: by assuming that the observed short-duration GRBs (931103B and 910711) with high-frequency QPO are originated from GFs of magnetar with torsional oscillation, the selected six EOSs (i.e., TM1, NL3, APR, SLy4, DDME2, and GM1) of NS which are suitable for the condition of zero-temperature (i.e., cold crust) exhibit significant overlapping in mass ranges, and it suggests that we can not constrain the EOS of NS via the high-frequency QPO in those two short-duration GBRs.
  • $\bullet$
@@
    Radial oscillation: by assuming that the observed short-duration GRBs (931103B and 910711) with high-frequency QPO are originated from the radial oscillation of HMNS from binary NS merger, the selected six EOSs (I.e., IUF, TM1, TMA, FSG, BHBLp, and NL3) of NS which are suitable for the condition of high-temperature (i.e., hot remnant) can not be ruled out with redshift considered. However, it is found that the different EOS of NS corresponds to different temperatures for a given redshift with $\text{Metric}=0$. It means that the EOS can only be constrained if the redshift and temperature of remnant can be measured.
@@
On the other hand, the origin of these two short GRBs may be related to the inspiral phase of the binary NS merger, which could also explain the observed high-frequency QPOs as arising from torsional oscillation. If this is the case, the $\gamma-$ray emission of short GRBs could result from the interaction of the NS magnetospheres during the inspiral phase (Vietri, 1996; Hansen & Lyutikov, 2001; McWilliams & Levin, 2011; Palenzuela et al., 2013; Wang et al., 2018), or from the cracking of the tidal crust in the last moments of a compact binary merger (Troja et al., 2010). In this framework, the GRB could be identified as an isolated precursor rather than being associated with relativistic outflows from post-merger remnants (Burlon et al., 2008; Zhong et al., 2019; Coppin et al., 2020; Wang et al., 2020). Unlike the fluid-dominated, high-temperature environment of an HMNS formed post-merger, the inspiral phase involves relatively cold NSs with intact crusts, where shear stresses can excite torsional modes. The maximum luminosity of the precursor from the interaction of NS magnetospheres is proportional to $B^{2}$. In fact, this requires introducing an additional assumption, namely, the binary system must contain at least one NS with a magnetar-like magnetic field ($B>10^{15}$ G) to achieve the observed luminosity (Wang et al., 2020).
@@
Recently, numerical relativity studies have suggested that the quasi-periodic kilohertz substructure observed in electromagnetic outflows from HMNSs may be dominated by magnetohydrodynamic (MHD) shearing processes rather than being related to internal stellar oscillations (Most & Quataert, 2023). This indicates that QPOs in relativistic outflows may not be necessarily related to internal NS oscillations, making it difficult to constrain the EOS of NS. This is consistent with the results obtained in this work. However, asteroseismology remains a potential method to constrain the EOS of NS when more precise information of the NS temperature can be obtained, such as the redshift and magnetic field strength. Moreover, Guedes et al. (2024) adopted numerical relativity simulations of post-merger remnants and their relationship to constrain the binary tidal coupling constant, and then constrain the EOS of NS by invoking the observed high-frequency QPO in those two short GRBs. They claim to have found tight constraints on the mass-radius relation, which differs significantly from what we have obtained in this paper. The possible reason is that the adopted method and the physical conditions are different from each other. Our method can constrain the redshift based on the presence of two distinct QPO frequencies ($\nu_{1}$ and $\nu_{2}$) in these bursts, with certain strong assumptions (e.g. temperature). In contrast, the approach from Guedes et al. (2024) adopted $\nu_{2}$ and the ratio $\nu_{2}/\nu_{1}$ to constrain the redshift, and the ratio between $\nu_{2}$ and $\nu_{1}$ is redshift-independent. We hope that multi-messenger observations, combining electromagnetic signals and gravitational waves, will offer more precise constraints on the physical properties of NS remnants in the future, such as their temperature, redshift, and magnetic field strength.

### v2

@@ 3 Radial Oscillation
The first direct evidence of short GRB origin, which comes from an NS–NS merger, is the detection of GW170817 by Advanced LIGO and Virgo associated with short GRB GRB 170817A and an optical/infrared transient known as kilonova AT 2017gfo (Abbott et al., 2017; Goldstein et al., 2017; Zhang et al., 2018). From a theoretical point of view, there are four different types of remnants of binary NS mergers, such as a BH, an HMNS, a supramassive NS supported by rigid rotation with a lifetime of hundreds of seconds before collapsing into a BH, and a stable NS (Dai & Lu, 1998a, b; Dai et al., 2006; Rosswog et al., 2000; Fan & Xu, 2006; Rezzolla et al., 2010; Giacomazzo & Perna, 2013; Zhang, 2013; Lasky et al., 2014; Lü et al., 2017; Gao et al., 2017). If the remnant is a newborn NS, the temperature is expected to reach several MeV, or even exceed tens of MeV in the immediate aftermath of the merger (Oechslin et al., 2007; Bauswein et al., 2010). Such high temperatures already exceed the point at which a crystalline lattice is stable, and the initial state of the remnant from NS–NS should be fluid filled. The cooling timescale is expected to range from thousands of seconds to days before the crystal forms. Therefore, torsional oscillations, which are shear forces from a crystalline crust, are unlikely to occur, but radial modes are more suitable (Stergioulas et al., 2011). The radial oscillation is driven by pressure and gravity, producing a radial displacement that is directly affected by the density and pressure profiles of NSs (Chandrasekhar, 1964; Chanmugam, 1977; Gondek et al., 1997a). Gondek et al. (1997b) found that the frequency spectrum of the lowest modes differs significantly from that of cold NSs. In this section, we will discuss the radial oscillation of NSs and constrain the EOS of NSs by adopting the high-frequency QPO of short-duration GRBs GRB 931103B and GRB 910711, assuming that these two short GRBs are indeed from an NS–NS merger.
@@ 3.1 Physical Model
Within the framework of general relativity, the study of infinitesimal radial adiabatic oscillations in stars was initially proposed by Chandrasekhar (1964) and applied to enhance its utility for numerical calculation by Chanmugam (1977) later. The study focused on two essential quantities, one being the radial displacement ($\Delta r$), which represents the radial motion of a matter element relative to its equilibrium position, and the other being the Lagrangian pressure perturbation ($\Delta P$), which quantifies the change in pressure associated with the displacement. These quantities are determined by a pair of ordinary differential equations, which are expressed as follows:
@@
Here $\xi=\Delta r/r$ is the normalized radial displacement. To avoid the singularity in $\xi$ at the stellar center and ensure a wellbehaved normalization for the eigenfunctions, we adopt $\xi(0)=1$ as the boundary condition. Parameter $\omega_{n}=2\pi\nu_{n}$ is the angular frequency of radial oscillation with the eigenfrequency $\nu_{n}$; the stability condition of oscillations requires $\omega^{2}\geq 0$. $\Gamma$ is the adiabatic index that is related to the baryon number density $n_{b}$ and pressure (Baym et al., 1971), namely,
@@
By solving Equations (8) and (9), it is necessary to apply boundary conditions at the stellar center and surface, namely,
@@
If this is the case, Equations (8) and (9) can be reduced to a second-order differential equation for $\xi(r)$ and switched to the form of a Sturm–Liouville eigenvalue problem. Similar to the case of torsional oscillation, the eigenvalues $\omega^{2}$ can also yield a discrete spectrum $\omega_{0}^{2}<\omega_{1}^{2}<\dots<\omega_{n}^{2}$, and each eigenvalue corresponds to an eigenfunction $\xi_{n}$ that has $n$ nodes within $0\leq r\leq R$. The fundamental mode ($n=0$), also known as the $f$-mode, corresponds to the lowest eigenvalue, while the higher modes with $n=1,2,3,\dots$ are referred to as $p$-modes (Sagun et al., 2020). The eigenfrequencies not only describe the nature of stellar oscillations but also serve as critical indicators of the system’s stability under small perturbations. Hence, all eigenvalues should satisfy $\omega_{n}^{2}\geq 0$ for every $n$ and ensure that a static stellar model is dynamically stable under small, radial, and adiabatic perturbations.
@@ 3.2 Constraint on the EOS of NS
An HMNS, supported by differential rotation that survives 10–100 ms before collapsing into a BH, is a possible remnant of a binary NS merger. Initially, the HMNS is expected to have an extremely high temperature, potentially reaching tens of MeV. Such high temperatures significantly affect the physical properties of HMNSs, such as the EOS of NSs. By assuming that the high-frequency QPOs in the short-duration GRBs GRB 931101B and GRB 910711 are caused by the radial oscillation of an HMNS from a binary NS merger, we employ six EOSs for NSs (NL3, IUF, TM1, TMA, FSG, and BHBLp), which are suitable for the condition of high temperature. The interpolated data for the selected EOS models were obtained from the publicly available CompOSE database [^1] https://compose.obspm.fr/table. (Oertel et al., 2017; Typel et al., 2015).
Figures 3 and 4 show the eigenfrequencies ($\nu_{n}$) of various EOS models as functions of the central density ($\rm log\rho_{c}$) with different temperatures for GRB 931101B and GRB 910711, respectively. The maximum temperature of five EOSs (i.e., IUF, TM1, TMA, FSG, and BHBLp) can reach 40 MeV, while for NL3 it extends up to as high as 60 MeV. It is found that the frequencies of $n=1$ modes at zero temperature are significantly higher than those of observed values for those two short GRBs. However, the frequency of the $n=1$ mode is rapidly decreased when the temperature is increased. The shaded vertical regions in Figures 3 and 4 correspond to the ranges of $\rho_{c}$ where the eigenfrequencies of $n=0$ and $n=1$ modes match the observed lower and higher frequencies of the QPO, respectively. If this is the case, the temperature at which the observed lower and higher frequencies of the QPO correspond to the same $\rho_{c}$ needs to be determined and can be used to constrain the EOS of NSs.
To find the temperature at which the vertical bands overlap, we define a metric: $\text{Metric}=\rho_{\rm gap}-\rho_{\rm cross}$, which quantifies the spatial relationship between the two vertical bands. Here the $\rho_{\rm gap}$ represents the distance between the closest edges of the two regions, namely,
@@
Here $[x_{0,\text{min}},x_{0,\text{max}}]$ and $[x_{1,\text{min}},x_{1,\text{max}}]$ are the $\rho_{c}$ ranges obtained by comparing the $n=0$ mode with the observed lower-frequency QPO and the $n=1$ mode with the higherfrequency QPO, respectively. If the bands overlap, $\rho_{\rm gap}=0$. Parameter $\rho_{\rm cross}$ is defined as
@@
If the bands do not overlap, $\rho_{\rm cross}<0$; otherwise, it is the size of the overlapping region between the two bands. $\text{Metric}=0$ implies that there exists a temperature at which the observed lower and higher frequencies of the QPO correspond to the same $\rho_{c}$ at a specific temperature. Figure 5 shows the relationship between the Metric and temperature ($T$) for different EOSs. It is found that most EOSs allow the Metric to reach a negative value at a specific temperature for those two short GRBs, except for the stiffest EOS, NL3. Hence, except NL3, we cannot rule out the other five EOSs based on the observed lower and higher frequencies of the QPO for those two short GRBs.
@@
On the other hand, the effect of redshift in the above calculation is ignored due to the lack of redshift measurements. In fact, one needs to consider the uncertainty caused by the redshift of these two short GRBs. Figure 6 shows the relationship between the temperature $T$ and the redshift $z$ for different EOSs. Here temperature $T$ values correspond to the points where the Metric reaches its minimum value. For example, the minimum value of temperature in Figure 5 for a given EOS corresponds to $z=0$. Then, we perform the calculations by varying the redshift range from 0 to 1 based on the observations. [^2] Lü et al. (2015) found that the redshift distribution for short GRBs with redshift measured is from 0 to 1, and average redshift is about $z=0.58$. We find that the temperature decreases as the redshift increases, and it is due to the relativistic effect of redshift that increases the frequency in the rest frame. Moreover, it is found that the NL3 EOS can reach the overlap of $\rho_{c}$ ranges at a higher temperature compared to other EOSs when $z\gtrsim 0.5$. For example, Figure 7 shows the eigenfrequency as a function of the NS mass for $n=0$ (solid lines) and $n=1$ (dashed lines) oscillation modes for different EOSs at $z=0$ and $z=1$, respectively.
On the other hand, the effect of redshift in the above calculation is ignored due to non-redshift measured. In fact, one needs to consider the uncertainty caused by the redshift of these two short GRBs. Fig. 6 shows the relationship between the temperature $T$ and the redshift $z$ for different EOSs. Here, temperature $T$ values are corresponding to the points where the Metric reaches its minimum value. For example, the minimum value of temperature in Fig. 5 for a given EOS is corresponding to $z=0$. Then, we perform the calculations by varying the redshift range from 0 to 1 based on the observations [^3] Lü et al. (2015) found that the redshift distribution for short GRBs with redshift measured is from 0 to 1, and the average redshift is about $z=0.58$.. We find that the temperature decreases as the redshift increases, and it is due to the relativistic effect of redshift which increases the frequency in the rest frame. Moreover, it is found that the NL3 EOS can reach the overlap of $\rho_{c}$ ranges at a higher temperature compared to other EOSs when $z\gtrsim 0.5$. For example, Fig. 7 shows the eigen-frequency as a function of the NS mass for $n=0$ (solid lines) and $n=1$ (dashed lines) oscillation modes for different EOSs at $z=0$ and $z=1$, respectively.
@@
In any case, it is difficult to constrain the EOS of NSs based solely on the observed high-frequency QPO in those two short GRBs, and we cannot rule out any EOS by considering the effect from both temperature and redshift. However, it is found that the different EOSs of NSs correspond to different temperatures for a given redshift with $\text{Metric}=0$. In other words, if we believe that the observed high-frequency QPOs in short GRBs are indeed caused by radial oscillations of HMNSs from binary NS mergers, one can constrain the EOS only when the redshift and temperature of the remnant can be measured.
@@ 4 Conclusion and Discussion
NSs offer unparalleled opportunities for investigating extreme physics phenomena in the universe. Gaining insights into the composition of matter and the structure of NSs would be a significant advancement in nuclear physics and astrophysics. The application of seismology in QPOs of SGRs has demonstrated immense potential in this regard, particularly in providing constraints on the EOS and the thickness of the crust of NSs. The recent identification of QPOs from two short-duration GRBs (GRB 931101B and GRB 910711) may provide a new opportunity to explore the nature of NSs by employing seismology.
@@
In this paper, we investigate the potential of torsional and radial oscillation models to explain the observed high-frequency QPOs in two short-duration GRBs (GRB 931101B and GRB 910711), and then we try to constrain the EOS of NSs. The following interesting results can be obtained:
  • $1$
@@
    Torsional oscillation. By assuming that the observed short-duration GRBs (GRB 931103B and GRB 910711) with high-frequency QPOs originated from GFs of magnetars with torsional oscillation, the selected six EOSs (i.e., TM1, NL3, APR, SLy4, DDME2, and GM1) of NSs that are suitable for the condition of zero temperature (i.e., cold crust) exhibit significant overlapping in mass ranges, and this suggests that we cannot constrain the EOS of NSs via the high-frequency QPOs in those two short-duration GBRs.
  • $2$
@@
    Radial oscillation. By assuming that the observed shortduration GRBs (GRB 931103B and GRB 910711) with high-frequency QPOs originated from the radial oscillation of an HMNS from a binary NS merger, the selected six EOSs (i.e., IUF, TM1, TMA, FSG, BHBLp, and NL3) of NSs that are suitable for the condition of high temperature (i.e., hot remnant) cannot be ruled out with redshift considered. However, it is found that the different EOSs of NSs correspond to different temperatures for a given redshift with Metric = 0. This means that the EOS can only be constrained if the redshift and temperature of the remnant can be measured.
@@
On the other hand, the origin of these two short GRBs may be related to the inspiral phase of the binary NS merger. This phase could also explain the observed high-frequency QPOs, which caused by torsional oscillation. If this is the case, the $\gamma$-ray emission of short GRBs could result from the interaction of the NS magnetospheres during the inspiral phase (Vietri, 1996; Hansen & Lyutikov, 2001; McWilliams & Levin, 2011; Palenzuela et al., 2013; Wang et al., 2018), or from the cracking of the tidal crust in the last moments of a compact binary merger (Troja et al., 2010). In this framework, the GRB could be identified as an isolated precursor rather than being associated with relativistic outflows from post-merger remnants (Burlon et al., 2008; Zhong et al., 2019; Coppin et al., 2020; Wang et al., 2020). Unlike the fluid-dominated, high-temperature environment of an HMNS formed post-merger, the inspiral phase involves relatively cold NSs with intact crusts, where shear stresses can excite torsional modes. The maximum luminosity of the precursor from the interaction of NS magnetospheres is proportional to $B^{2}$ In fact, this requires introducing an additional assumption, namely, the binary system must contain at least one NS with a magnetar-like magnetic field ($B>10^{15}$ G) to achieve the observed luminosity (Wang et al., 2020).
@@
Recently, numerical relativity studies have suggested that the quasi-periodic kHz substructure observed in electromagnetic outflows from HMNSs may be dominated by magnetohydro-dynamic (MHD) shearing processes rather than being related to internal stellar oscillations (Most & Quataert, 2023). This indicates that QPOs in relativistic outflows may not be necessarily related to internal NS oscillations, making it difficult to constrain the EOS of NSs. This is consistent with the results obtained in this work. However, asteroseismology remains a potential method to constrain the EOS of NSs when more precise information of the NS temperature can be obtained, such as the redshift and magnetic field strength. Moreover, Guedes et al. (2024) adopted numerical relativity simulations of post-merger remnants and their relationship to constrain the binary tidal coupling constant, and then they constrain the EOS of NSs by invoking the observed highfrequency QPOs in those two short GRBs. They claim to have found tight constraints on the mass–radius relation, which differs significantly from what we have obtained in this paper. The possible reason is that the adopted method and the physical conditions are different from each other. Our method can constrain the redshift based on the presence of two distinct QPO frequencies ($\nu_{1}$ and $\nu_{2}$) in these bursts, with certain strong assumptions (e.g., temperature). In contrast, the approach from Guedes et al. (2024) adopted $\nu_{2}$ and the ratio $\nu_{2}/\nu_{1}$ to constrain the redshift, and the ratio between $\nu_{2}$ and $\nu_{1}$ is redshift independent. We hope that multimessenger observations, combining electromagnetic signals and GWs, will offer more precise constraints on the physical properties of NS remnants in the future, such as their temperature, redshift, and magnetic field strength.